OpenStudy (anonymous):

y varies directly as the square of x. When x = 1, y = 3. What is the value of y when x = 3? 1 9 27 81

OpenStudy (accessdenied):

if its a direct variation, what does the equation relating x and y look like?

OpenStudy (anonymous):

do i add both x & y ? idk how to do it

OpenStudy (anonymous):

y = kx^2 3 = k*1 so y = 3x^2 when x = 3 y = 27

OpenStudy (accessdenied):

well, in general, a direct variation looks like: y = kx In this case, we have "y varies directly as the SQUARE of x" so, y = kx^2 would suit this problem nicely

OpenStudy (accessdenied):

..yeah, that's correct. :P